# DynaSolidGeo: A Dynamic Benchmark for Genuine Spatial Mathematical Reasoning of VLMs in Solid Geometry

### v1

@@ 1 Introduction
Geometry problem solving has long played a central role in mathematical reasoning, requiring integrating visual understanding and symbolic reasoning across complex graphic and textual contexts [38]. According to structural properties, geometry can be categorized into plane geometry and solid geometry. Compared to plane geometry, solid geometry imposes substantially higher spatial mathematical reasoning ability, as reasoning in three dimensions entails spatial intelligence, including spatial perception, spatial relation, spatial orientation, spatial rotation, and spatial visualization that goes beyond two-dimensional recognition [14, 29, 16, 22]. Such tasks remain difficult even for human learners [7], and represent a formidable open challenge for current AI systems.
@@
In parallel, recent years have witnessed remarkable progress in multimodal large language models (MLLMs). Building on the successes of foundation models, vision-language models (VLMs) [23, 6, 11, 27, 5, 4, 17] have rapidly advanced the state of the art in a wide spectrum of multimodal understanding tasks. Among these tasks, multimodal mathematical reasoning has emerged as a challenging yet vibrant frontier, with benchmarks such as GeoQA [8], MathVista [25], and GeoSense [38] exposing both the promise and the limitations of current VLMs. These carefully designed benchmarks have played a pivotal role in advancing the field, providing standardized evaluation and catalyzing iterative improvements in model design and training paradigm.
@@
Despite this progress, current multimodal mathematical reasoning benchmarks exhibit critical limitations. First, the vast majority of existing geometry-related benchmarks focus on plane geometry or diagram-based word problems, leaving solid geometry, which places heightened demands for spatial intelligence, largely underexplored. For example, PGPS9K [42] contains more than 9,000 plane geometry questions but no solid geometry items, and in GeoEval [41], tasks involving solid geometry constitute merely 2% of the benchmark. Second, nearly all existing multimodal mathematical reasoning benchmarks are static, relying on fixed and finite test sets that are susceptible to data contamination and memorization. Recent analyses demonstrate that large models can memorize and regurgitate benchmark data [26, 12, 21, 28, 10], and some studies show that decontaminated re-releases often lead to substantial drops in performance [44], confirming that static evaluation may significantly overestimate true reasoning and generalization [31]. Similar concerns have motivated dynamic evaluations in coding [18, 45] and general-purpose QA [34, 35], indicating a community-wide shift towards dynamic, contamination-resistant evaluation protocols [9]. Moreover, most existing multimodal mathematical reasoning benchmarks [25, 33, 32, 41, 40, 42, 47] evaluate models solely based on answer accuracy, which allows models suffering from data contamination or over-reliance on memorization to appear strong, while failing to reveal their genuine reasoning ability.
@@
To address these limitations, we present DynaSolidGeo, a new benchmark for the dynamic evaluation of VLMs’ genuine spatial mathematical reasoning in solid geometry. Unlike existing static resources, DynaSolidGeo consists of 503 seed questions of solid geometry problem solving, each represented by a Python program paired with a corresponding MATLAB program. With the correctness of the question guaranteed, each seed question is parameterized: textual variables in the question statement (e.g., endpoint labels, side lengths, areas, volumes, ratio) as well as rendering parameters of the solid geometry (e.g., camera viewpoints) can all be randomized. By supplying different random seeds, DynaSolidGeo can in principle generate an unbounded number of question-answer instances, where each instance can optionally include two visual versions: a randomized-view image and a 360-degree rotation video. The seed questions of DynaSolidGeo are drawn from diverse and authoritative sources, including China’s Gaokao examinations, international mathematics competitions, and widely used training materials for competition preparation. Together, they cover nearly all major categories of high-school and competition-level solid geometry problems (eight in total), including positional relations, angle, length, area, and volume calculations, as well as counting, dynamic, and folding tasks. Moreover, we move beyond answer-only evaluation by incorporating process-level assessment grounded in expert-annotated reasoning chains. Through Answer Accuracy (AA), Process Score (PS), and Process-Qualified Accuracy (PA), we jointly measure answer correctness, reasoning quality, and reasoning-qualified accuracy, offering a more faithful reflection of VLMs’ genuine spatial mathematical reasoning ability. To ensure reliability, all solutions are expert-annotated by undergraduates and graduate students from the School of Mathematical Sciences, Peking University, including Chinese Mathematical Olympiad (CMO) gold medalists. A comparative summary with related benchmarks is provided in Table 1.
@@ 2 Related Work
Recent years have witnessed the emergence of multimodal benchmarks that evaluate mathematical reasoning in visually grounded settings. Early efforts include TQA [19] and Geometry3K [24] introduced multimodal reasoning tasks involving diagram-based science and geometry word problems with accompanying 2D visuals. More recent benchmarks, such as GeoQA [8], PGPS9K [42], MMMU-MATH [40], GeoEval [41], MATH-Vision [32], OlympiadBench [15], MathVerse [43], GeoSense [38], and GeoLaux [13] have broadened coverage to thousands of multimodal math problems. However, the vast majority of these resources concentrate on plane geometry and 2D diagrammatic reasoning, leaving solid geometry largely underexplored. A few datasets have attempted to move toward 3D: SolidGeo [33] explicitly targets solid geometry but remains static datasets vulnerable to contamination and memorization; DynaMath [47] introduces dynamic instance generation, but solid geometry is barely represented, with only 15 problems (3%) in the dataset. In contrast, DynaSolidGeo fills this gap with scalable and dynamic solid-geometry coverage.
@@
Recent vision–language models (VLMs) such as BLIP-2 [20], Flamingo [3], and LLaVA [23] combine pretrained large language models with visual encoders, enabling open-ended multimodal reasoning and instruction following. Building on this paradigm, the latest generation of VLMs has rapidly advanced in scale, architecture, and reasoning capability. The closed-source models include GPT-5 family [27], Gemini-2.5 family [11], and Claude-Sonnet-4.5 [5], which feature deeply integrated multimodal backbones and enhanced reasoning modules. In parallel, the open-source community has introduced competitive alternatives such as LLaVA-OneVision-1.5 family [4], GLM-4.1V-9B-Thinking [17], Llama-4-Maverick-17B-Instruct [2], InternVL3.5-8B [37], DeepSeek-VL2 [36], and the Qwen3-VL family [30], which push the frontier of visual capabilities. Yet their spatial mathematical reasoning ability remains underexplored, motivating our evaluation on DynaSolidGeo.
@@ 3.2 Data Annotation Pipeline
Expert-Guided Parametrization and Visualization: Here we follows a human-in-the-loop strategy, where human experts collaborate with large models to create a JSON annotation and a MATLAB program for each source question:
JSON-Based Question Parametrization: For each collected source question, mathematics experts parameterize the question statement by converting fixed values into variable parameters (e.g., endpoint labels, side lengths, areas, volumes, ratios) using f-string syntax, while ensuring correctness and availability. Corresponding answer is also expressed in terms of these variable parameters. An example of parameterized variables is highlighted in red in Fig. 1. Additional metadata (e.g., category and difficulty level) is also included and stored in JSON format.
@@ 3.3 Statistics
Table 2 summarizes the detailed statistics of the DynaSolidGeo dataset. In total, the benchmark contains 503 curated seed questions, all newly constructed for this work. The questions span a diverse set of solid geometry problem categories, including Positional relationship determination (PD, 11.7%), Angle calculation (AN, 20.5%), Length and distance calculation (LC, 13.1%), Area calculation (AR, 11.3%), Volume calculation (VC, 10.3%), Counting problems (CP, 7.0%), Dynamic or moving-point problems (DM, 13.1%), and Folding and unfolding problems (FP, 12.9%). The distribution across difficulty levels is reasonably balanced, with 27.2% easy, 57.7% medium, and 15.1% hard questions. Regarding question types, previous benchmarks (e.g., DynaMath[47], SolidGeo [33], GeoSense[38], MathVerse[43]) generally include multiple-choice questions. Such options inevitably provide strong hints to the models, thereby reducing the difficulty and making it difficult to assess their reasoning ability accurately. In contrast, we rewrite the multiple-choice and proof questions from the source data into fill-in-the-blank formats in our work. As a result, DynaSolidGeo consists of 88.3% numerical questions and 11.7% free-form questions, posing greater challenges to the reasoning ability of VLMs.
@@ 3.4.2 Process Score
To more accurately assess the quality of the model’s reasoning process, we annotate not only parameterized question statements and answers, but also parameterized reference reasoning processes. Each reasoning process is evaluated using LLM as a judge according to the following criteria:
@@ 4 Experiment
DynaSolidGeo supports the random generation of two visual versions: a randomized-view image and a 360-degree rotation video, for each question instance. Since existing geometry problem-solving tasks focus exclusively on the text-image modality, here we also evaluate models under the same text-image setting.
@@ 4.1 Experimental Setup
Implementation Details. We sample batches of question instances by setting the random seed to 0, 1, and 2, respectively, resulting in a total of 1,509 text-image question instances. For answer evaluation, we allow a 1% relative error tolerance. For process evaluation, we employ Qwen3-14B [39] as the judge model. For the evaluated models, we deploy small-scale models, including the LLaVA-OneVision-1.5 family, GLM-4.1V-9B-Thinking, InternVL3.5-8B, and the Qwen3-VL family (4B, 30B), on NVIDIA A800 GPUs for evaluation. DeepSeek-VL2 is evaluated via the SiliconFlow API111https://www.siliconflow.com/, while all remaining models are accessed through the OpenRouter API222https://openrouter.ai/ for evaluation.
@@ 4.2 Experimental Results
Performance Differences across Categories from the Perspective of Spatial Intelligence. As shown in Table 3, the best-performing models perform well in Area calculation (AR), Volume calculation (VC), and Dynamic or moving-point problems (DM). However, all models struggle with Counting problems (CP), and a significant performance gap can be observed between open-source and closed-source models in this category. This discrepancy can be explained through the lens of spatial intelligence theory [14]. Tasks such as AR, VC, and DM mainly rely on lower or mid-level spatial perception, spatial relation, and spatial orientation, where visual cues are explicit and reasoning can be simplified into formula- or rule-based deduction, rather than fine-grained 3D structural reasoning. This aligns well with the representational strengths of current MLLMs, which operate on visual encodings and symbolic reasoning. In contrast, Counting Problems (CP) require higher-order mental rotation and spatial visualization, requiring 3D reconstruction, occlusion reasoning, and mental manipulation of hidden or rotated objects. This explains why models perform relatively well on AR, VC, and DM tasks but fail consistently on CP tasks.
Metric Degradation under Process Evaluation. As shown in Table 3, after introducing process evaluation, all models exhibit varying degrees of decline in both PS and PA metrics compared to AA. Among them, Gemini-2.5-Pro shows the largest drop, with PA decreasing by 9.4% relative to AA, followed by Llama-3.2-90B-Vision-Instruct, whose PA drops by 6%. This suggests that these models, while capable of producing correct answers, often rely on reasoning processes that are less coherent or causally aligned with the final answers. Furthermore, the decline in PA relative to AA is generally smaller for thinking models than for instruct models. For example, GLM-4.1V-9B-Thinking shows only a 1.5% drop, and Qwen3-VL-8B-Thinking decreases by merely 0.1%, whereas both Llama-3.2-90B-Vision-Instruct and Llama-4-Maverick-17B-Instruct experience drops exceeding 5%. Even within the Qwen3-VL family, Qwen3-VL-8B-Thinking and Qwen3-VL-30B-A3B-Thinking exhibit smaller declines compared to Qwen3-VL-8B-Instruct, Qwen3-VL-30B-A3B-Instruct, and Qwen3-VL-235B-A22B-Instruct. These observations suggest that thinking models generally produce reasoning processes that are more coherent, logically sound, and causally consistent with their final answers than those of instruct models. In addition, only a few stronger models, such as those from the Gemini and Qwen3-VL family, achieve identical AA, PS, and PA on Counting problems (CP), whereas others show clear metric gaps on different task types. This indicates that the CP task requires higher-order spatial intelligence and more rigorous symbolic reasoning. Consequently, making correct answers less susceptible to hallucination or logical inconsistency.
@@ 4.3 Error Analysis
We conduct an error analysis on two representative models, Gemini-2.5-Pro (closed source) and Qwen3-VL-30B-A3B-Thinking (open source). Specifically, we categorize the errors in Process-Qualified Accuracy (PA) into five types: Visual Perception Errors, Logical Reasoning Errors, Calculation Errors, Knowledge Errors, and Hallucination Errors, as shown in Figure 4. Across the 1,509 sampled instances, Gemini-2.5-Pro makes a total of 715 errors, Qwen3-VL-30B-A3B-Thinking makes 525 errors. Among these error types, visual perception and logical reasoning errors dominate, with visual perception errors accounting for the largest proportion. This indicates that, although these advanced models have demonstrated strong symbolic reasoning capabilities, they still lack sufficient perceptual understanding in solid geometry tasks that require spatial intelligence. Qwen3-VL-30B-A3B-Thinking exhibits 52 more Visual Perception errors than Gemini-2.5-Pro, suggesting that it is more prone to reasoning failures triggered by inaccurate visual perception. This also explains why Qwen3-VL-30B-A3B-Thinking performs significantly worse than Gemini-2.5-Pro on Counting Problems (CP), which demand higher levels of spatial intelligence. In addition, Qwen3-VL-30B-A3B-Thinking makes fewer errors in other categories (including Logical Reasoning Errors), demonstrating its stronger symbolic reasoning capability.

### v2

@@ 1 Introduction
Geometry problem solving has long played a central role in mathematical reasoning, requiring integrating visual understanding and symbolic reasoning across complex graphic and textual contexts [48, 50, 42, 63, 35, 32, 44, 30, 43, 10, 21]. According to structural properties, geometry can be categorized into plane geometry and solid geometry. Compared to plane geometry, solid geometry imposes substantially higher spatial mathematical reasoning ability, as reasoning in three dimensions entails spatial intelligence, including spatial perception, spatial relation, spatial orientation, spatial rotation, and spatial visualization that goes beyond two-dimensional recognition [51, 53, 16, 36, 19, 26]. Such tasks remain difficult even for human learners [7], and represent a formidable open challenge for current AI systems [59].
@@
In parallel, recent years have witnessed remarkable progress in multimodal large language models (MLLMs). Building on the successes of foundation models, vision-language models (VLMs) [27, 6, 12, 33, 5, 4, 20, 14, 17, 60] have rapidly advanced the state of the art in a wide spectrum of multimodal understanding tasks. Among these tasks, multimodal mathematical reasoning has emerged as a challenging yet vibrant frontier, with benchmarks such as GeoQA [8], MathVista [29], GeomRel [42], and GeoSense [50] exposing both the promise and the limitations of current VLMs. These carefully designed benchmarks have played a pivotal role in advancing the field, providing standardized evaluation and catalyzing iterative improvements in model design and training paradigm.
@@
Despite this progress, current multimodal mathematical reasoning benchmarks exhibit critical limitations. First, the vast majority of existing geometry-related benchmarks focus on plane geometry or diagram-based word problems, leaving solid geometry, which places heightened demands for spatial intelligence, largely underexplored. For example, PGPS9K [57] contains more than 9,000 plane geometry questions but no solid geometry items, and in GeoEval [56], tasks involving solid geometry constitute merely 2% of the benchmark. Second, nearly all existing multimodal mathematical reasoning benchmarks are static, relying on fixed and finite test sets that are susceptible to data contamination and memorization. Recent analyses demonstrate that large models can memorize and regurgitate benchmark data [31, 13, 25, 34, 11, 37], and some studies show that decontaminated re-releases often lead to substantial drops in performance [62], confirming that static evaluation may significantly overestimate true reasoning and generalization [39]. Similar concerns have motivated dynamic evaluations in coding [22, 64] and general-purpose QA [45, 46], indicating a community-wide shift towards dynamic, contamination-resistant evaluation protocols [9, 61]. Moreover, most existing multimodal mathematical reasoning benchmarks [29, 41, 40, 56, 55, 57, 66] evaluate models solely based on answer accuracy, which allows models suffering from data contamination or over-reliance on memorization to appear strong, while failing to reveal their genuine reasoning ability.
@@
To address these limitations, we present DynaSolidGeo, a new benchmark for the dynamic evaluation of VLMs’ genuine spatial mathematical reasoning in solid geometry. Unlike existing static resources, DynaSolidGeo consists of 503 seed questions of solid geometry problem solving, each represented by a Python program paired with a corresponding MATLAB program. With the correctness of the question guaranteed, each seed question is parameterized: textual variables in the question statement (e.g., endpoint labels, side lengths, areas, volumes, ratios) as well as rendering parameters of the solid geometry (e.g., camera viewpoints) can all be randomized. By supplying different random seeds, DynaSolidGeo can, in principle, generate an unbounded number of question-answer instances, where each instance can optionally include two visual versions: a randomized-view image and a 360-degree rotation video. The seed questions of DynaSolidGeo are drawn from diverse and authoritative sources, including China’s Gaokao examinations, international mathematics competitions, and widely used training materials for competition preparation. Together, they cover nearly all major categories of high-school and competition-level solid geometry problems (eight in total), including positional relations, angle, length, area, and volume calculations, as well as counting, dynamic, and folding tasks. Moreover, we move beyond answer-only evaluation by incorporating process-level assessment grounded in expert-annotated reasoning chains. Through Answer Accuracy (AA), Process Score (PS), and Process-Qualified Accuracy (PA), we jointly measure answer correctness, reasoning quality, and reasoning-qualified accuracy, offering a more faithful reflection of VLMs’ genuine spatial mathematical reasoning ability. To ensure reliability, all solutions are expert-annotated by undergraduates and graduate students from the School of Mathematical Sciences, Peking University, including Chinese Mathematical Olympiad (CMO) gold medalists. A comparative summary with related benchmarks is provided in Table 1.
@@ 2 Related Work
Recent years have witnessed the emergence of multimodal benchmarks that evaluate mathematical reasoning in visually grounded settings. Early efforts include TQA [23] and Geometry3K [28] introduced multimodal reasoning tasks involving diagram-based science and geometry word problems with accompanying 2D visuals. More recent benchmarks, such as GeoQA [8], PGPS9K [57], MMMU-MATH [55], GeoEval [56], MATH-Vision [40], OlympiadBench [18], MathVerse [58], GeomRel [42], GeoSense [50], and GeoLaux [15] have broadened coverage to thousands of multimodal math problems. However, most of these resources focus on plane geometry and 2D diagrammatic reasoning, leaving solid geometry largely underexplored. A few datasets have attempted to move toward 3D: SolidGeo [41] explicitly targets solid geometry but remains static datasets vulnerable to contamination and memorization; DynaMath [66] introduces dynamic instance generation, but solid geometry is barely represented, with only 15 problems (3%) in the dataset. In contrast, DynaSolidGeo fills this gap with scalable and dynamic solid-geometry coverage.
@@
Recent vision–language models (VLMs) such as BLIP-2 [24], Flamingo [3], and LLaVA [27] combine pretrained large language models with visual encoders, enabling open-ended multimodal reasoning and instruction following. Building on this paradigm, the latest generation of VLMs has rapidly advanced in scale, architecture, and reasoning capability. The closed-source models include GPT-5 family [33], Gemini-2.5 family [12], and Claude-Sonnet-4.5 [5], which feature deeply integrated multimodal backbones and enhanced reasoning modules. In parallel, the open-source community has introduced competitive alternatives such as LLaVA-OneVision-1.5 family [4], GLM-4.1V-9B-Thinking [20], Llama-4-Maverick-17B-Instruct [2], InternVL3.5-8B [49], DeepSeek-VL2 [47], and the Qwen3-VL family [38], which push the frontier of visual capabilities. Yet their spatial mathematical reasoning ability remains underexplored, motivating our evaluation of DynaSolidGeo.
@@ 3.2 Data Annotation Pipeline
Expert-Guided Parametrization and Visualization: Here we follow a human-in-the-loop strategy, where human experts collaborate with large models to create a JSON annotation and a MATLAB program for each source question:
JSON-Based Question Parametrization: For each collected source question, mathematics experts parameterize the question statement by converting fixed values into variable parameters (e.g., endpoint labels, side lengths, areas, volumes, ratios) using f-string syntax, while ensuring correctness and availability. The corresponding answer is also expressed in terms of these variable parameters. An example of parameterized variables is highlighted in red in Fig. 1. Additional metadata (e.g., category and difficulty level) is also included and stored in JSON format.
@@ 3.3 Statistics
Table 2 summarizes the detailed statistics of the DynaSolidGeo dataset. In total, the benchmark contains 503 curated seed questions, all newly constructed for this work, and provides both Chinese and English versions for each question statement. The questions span a diverse set of solid geometry problem categories, including Positional relationship determination (PD, 11.7%), Angle calculation (AN, 20.5%), Length and distance calculation (LC, 13.1%), Area calculation (AR, 11.3%), Volume calculation (VC, 10.3%), Counting problems (CP, 7.0%), Dynamic or moving-point problems (DM, 13.1%), and Folding and unfolding problems (FP, 12.9%). The distribution across difficulty levels is reasonably balanced, with 27.2% easy, 57.7% medium, and 15.1% hard questions. Regarding question types, previous benchmarks (e.g., DynaMath[66], SolidGeo [41], GeoSense[50], MathVerse[58]) generally include multiple-choice questions. Such options inevitably provide strong hints to the models, thereby reducing the difficulty and making it difficult to assess their reasoning ability accurately. In contrast, we rewrite the multiple-choice and proof questions from the source data into fill-in-the-blank formats in our work. As a result, DynaSolidGeo consists of 88.3% numerical questions and 11.7% free-form questions, posing greater challenges to the reasoning ability of VLMs.
@@ 3.4.2 Process Score
To more accurately assess the quality of the reasoning process, we also annotate parameterized reference reasoning chains. Based on these expert-annotated reasoning chains, each reasoning process is evaluated using LLM as a judge according to the following criteria:
@@ 4 Experiment
DynaSolidGeo supports the random generation of two visual versions: a randomized-view image and a 360-degree rotation video, for each question instance. Since existing geometry problem-solving tasks focus exclusively on the text-image modality, here we also evaluate models under the same text-image setting. The performance of some VLMs in the text-video modality can be found in Appendix D.2.
@@ 4.1 Experimental Setup
Implementation Details. We sample batches of question instances by setting the random seed to 0, 1, and 2, respectively, resulting in a total of 1,509 text-image question instances. For answer evaluation, we allow a 1% relative error tolerance. For process evaluation, we employ Qwen3-14B [52] as the judge model. For the evaluated models, we deploy small-scale models, including the LLaVA-OneVision-1.5 family, GLM-4.1V-9B-Thinking, InternVL3.5-8B, and the Qwen3-VL family (4B, 30B), on NVIDIA A800 GPUs for evaluation. DeepSeek-VL2 is evaluated via the SiliconFlow API111https://www.siliconflow.com/, while all remaining models are accessed through the OpenRouter API222https://openrouter.ai/ for evaluation. We set the temperature to 0.0 for all models to reduce randomness, while keeping all other hyperparameters at their default values. More details of the experiment setup can be found in Appendix C.
@@ 4.2 Experimental Results
Performance Differences across Categories from the Perspective of Spatial Intelligence. As shown in Table 3, the best-performing models perform well in Area calculation (AR), Volume calculation (VC), and Dynamic or moving-point problems (DM). However, all models struggle with Counting problems (CP), and a significant performance gap can be observed between open-source and closed-source models in this category. This discrepancy can be explained through the lens of spatial intelligence theory [16] (for more background on spatial intelligence, see Appendix B). Tasks such as AR, VC, and DM mainly rely on lower or mid-level spatial perception, spatial relation, and spatial orientation, where visual cues are explicit and reasoning can be simplified into formula- or rule-based deduction, rather than fine-grained 3D structural reasoning. This aligns well with the representational strengths of current MLLMs, which operate on visual encodings and symbolic reasoning. In contrast, Counting Problems (CP) require higher-order mental rotation and spatial visualization, requiring 3D reconstruction, occlusion reasoning, and mental manipulation of hidden or rotated objects. This explains why models perform relatively well on AR, VC, and DM tasks but fail consistently on CP tasks.
Metric Degradation under Process Evaluation. As shown in Table 3, after introducing process evaluation, all models exhibit varying degrees of decline in both PS and PA metrics compared to AA. Among them, Gemini-2.5-Pro shows the largest drop, with PA decreasing by 9.4% relative to AA, followed by Llama-3.2-90B-Vision-Instruct, whose PA drops by 6%. This suggests that these models, while capable of producing correct answers, often rely on reasoning processes that are less coherent or causally aligned with the final answers. Furthermore, the decline in PA relative to AA is generally smaller for thinking models than for instruct models. For example, GLM-4.1V-9B-Thinking shows only a 1.5% drop, and Qwen3-VL-8B-Thinking decreases by merely 0.1%, whereas both Llama-3.2-90B-Vision-Instruct and Llama-4-Maverick-17B-Instruct experience drops exceeding 5%. Even within the Qwen3-VL family, Qwen3-VL-8B-Thinking and Qwen3-VL-30B-A3B-Thinking exhibit smaller declines compared to Qwen3-VL-8B-Instruct, Qwen3-VL-30B-A3B-Instruct, and Qwen3-VL-235B-A22B-Instruct. These observations suggest that thinking models generally produce reasoning processes that are more coherent, logically sound, and causally consistent with their final answers than those of instruct models. In addition, only a few stronger models, such as those from the Gemini and Qwen3-VL family, achieve identical AA, PS, and PA on Counting problems (CP), whereas others show clear metric gaps on different task types. This indicates that the CP task requires higher-order spatial intelligence and more rigorous symbolic reasoning, consequently making correct answers less susceptible to hallucination or logical inconsistency.
@@ 4.3 Error Analysis
We conduct an error analysis on two representative models, Gemini-2.5-Pro (closed source) and Qwen3-VL-30B-A3B-Thinking (open source). Specifically, we categorize the errors in Process-Qualified Accuracy (PA) into five types: Visual Perception Errors, Logical Reasoning Errors, Calculation Errors, Knowledge Errors, and Hallucination Errors, as shown in Figure 4. Across the 1,509 sampled instances, Gemini-2.5-Pro makes a total of 715 errors, Qwen3-VL-30B-A3B-Thinking makes 525 errors. Among these error types, visual perception and logical reasoning errors dominate, with visual perception errors accounting for the largest proportion. This indicates that, although these advanced models have demonstrated strong symbolic reasoning capabilities, they still lack sufficient perceptual understanding in solid geometry tasks that require spatial intelligence. Qwen3-VL-30B-A3B-Thinking exhibits 52 more Visual Perception errors than Gemini-2.5-Pro, suggesting that it is more prone to reasoning failures triggered by inaccurate visual perception. This also explains why Qwen3-VL-30B-A3B-Thinking performs significantly worse than Gemini-2.5-Pro on Counting Problems (CP), which demand higher levels of spatial intelligence. In addition, Qwen3-VL-30B-A3B-Thinking makes fewer errors in other categories (including Logical Reasoning Errors), demonstrating its stronger symbolic reasoning capability. More details on the error analysis can be found in Appendix D.3.